Hello,
I'm trying to find the number of years for an equation using compound interest.
The question is;
How many years will it take for a $2400 sapphire ring to be worth $8000 if it increases in value by 10.6% pa? Round answer to nearest whole number in years.
I've identified
P= 2400
A= 8000
r= 10.6

Answer by mananth(16946)

Find number of years
Principal P = 2400
Amount= 8000
n= number of periods 4
compounded (t) 1
Rate = 10.6 0
n = ln(FV / PV) / ln(1 + r) (r/q))

n=( ln ( 8000 / 2400 )/ln( 1 + 0.11 )
n=( 1.2 )/ ln 1.11
n=( 1.2 / 0.1 )
n= 11.95 quarters
11.95 years

12 years